A company is selling an equipment after four years for $37,249. The equipment was originally purchased for $88,575. The tax rate is 12%.The equipment is classified as a 5-year property. What is the after-tax salvage value?

The MACRS allowance percentages are as follows, starting with Year 1: 20.00, 32.00, 19.20, 11.52, 11.52, and 5.76 percent.

Enter your answer rounded off to two decimal points. Do not enter $ or comma in the answer box.
Answer #1

After tax salvage value is $ 33,551.52

Step-1:Calculation of book value after 4 years
Year Cost Depreciation rate Depreciation Expense Accumulated depreciation Book Value
a b c=a*b d e=a-d
1 $ 37,249.00 20.00% $      7,449.80 $      7,449.80 $ 29,799.20
2 $ 37,249.00 32.00% $    11,919.68 $   19,369.48 $ 17,879.52
3 $ 37,249.00 19.20% $      7,151.81 $   26,521.29 $ 10,727.71
4 $ 37,249.00 11.52% $      4,291.08 $   30,812.37 $   6,436.63
Step-2:Calculation of after tax salvage value
Sales value a $    37,249.00
Book Value b $      6,436.63
Profit on sale c=a-b $    30,812.37
Tax on profit on sale d=c*12% $      3,697.48
After tax salvage value e=a-d $    33,551.52